# Robust Bernoulli mixture models for credit portfolio risk

### v1

@@ Acknowledgements
This appendix addresses the proofs of all results in this paper. We move the proof of Theorem 2.1 after the proof of Proposition 4.1 because it is based on several results from Sections 3 and 4,
Proof of Lemma 3.1:.
@@
$P(D_{n}=1\mid Z=z)$ $=P(X_{n}\leq c_{n}\mid Z=z)$ (5.12)
$=P(f_{n}(Y,\varepsilon_{n})\leq c_{n}\mid g(Y)=z)$ (5.13)
$=P(f_{n}(g^{-1}(z),\varepsilon_{n})\leq c_{n}\mid Y=g^{-1}(z))$ (5.14)
$=P(f_{n}(g^{-1}(z),\varepsilon_{n})\leq c_{n})\,,$ (5.15)
@@
Under the continuity assumptions, it can easily be verified that $X_{n}$ has a continuous distribution function. Hence, $c_{n}=F_{X_{n}}^{-1}(\pi_{n})$ implies $\pi_{n}=P(X_{n}\leq c_{n})=P(D_{n}=1)\,.$
Proof of Proposition 3.3:.
@@
$F_{X_{n},Y}(x,y)=P(X_{n}\leq x,Y\leq y)$ $=\int_{0}^{y}P(f_{n}(Y,\varepsilon_{n})\leq x\mid Y=v)\mathrm{\,d}v$ (5.16)
$=\int_{0}^{y}P(f_{n}(v,\varepsilon_{n})\leq x)\mathrm{\,d}v$ (5.17)
$=\int_{0}^{y}P(C_{n}^{-1}(\varepsilon_{n}|v)\leq x)\mathrm{\,d}v$ (5.18)
$=\int_{0}^{y}P(\varepsilon_{n}\leq C_{n}(x|v))\mathrm{\,d}v$ (5.19)
$=\int_{0}^{y}C_{n}(x|v)\mathrm{\,d}v$ (5.20)
$=\int_{0}^{y}\partial C_{n}(x,v)/\partial v\mathrm{\,d}v=C_{n}(x,y)=C_{X_{n},Y}(x,y)\,,$ (5.21)
where the second equality follows from independence of $Y$ and $\varepsilon_{n}\,.$ For the fifth equality, we use that $\varepsilon_{n}$ is uniformly distributed on $(0,1)\,.$
@@
To show statement (ii), first observe that $f_{n}(v,t)$ is increasing in its second component $t$ because $C_{n}(t|v)$ and thus $C_{n}^{-1}(t|v)$ are increasing in $t\,.$ Further, $f_{n}(v,t)=C_{n}^{-1}(t|v)$ is increasing in $v$ for all $t\,$. This follows from the fact that $C_{n}$ is SI, which in turn implies the concavity of $C_{n}(t,v)$ in $v$ for all $t$, and the fact that $C_{n}(t|v)=(\partial C_{n}(t,v)/\partial v)$ is decreasing in $v$ for all $t$ if and only if $C_{n}^{-1}(t|v)$ is increasing in $v$ for all $t$ (note that the inverse is w.r.t. the first component $t$ while the distribution functions are compared for different $v$).
Since $X_{n}\sim U(0,1)\,,$ the default indicator $D_{n}$ satisfies condition (I). Properties (II) and (III) follow from Lemma 3.1. ∎
Proof of Lemma 3.5:.
We construct $X_{n}$ and $Y$ through the distributional transform of $D_{n}$ and $Z\,.$ To this end, define for a random variable $\xi$ its distributional transform by
$\tau_{\xi}:=F_{\xi}(\xi,W):=F_{\xi}(\xi-)+W(F_{\xi}(\xi)-F_{\xi}(\xi-))\,,$ (5.22)
where $W\sim U(0,1)$ is independent from $\xi$ and where $F(x-)$ denotes the left-hand limit of $F$ at $x\,.$ It holds that
$\tau_{\xi}$ $\sim U(0,1)\,,$ (5.23)
[…]
@@
$X_{n}:=1-X_{n}^{\prime}\quad\text{and}\quad Y:=1-Y^{\prime}\,.$ (5.25)
Then (5.23) implies $X_{n},Y\sim U(0,1)\,.$ Further, we have
$X_{n}$ $\leq c_{n}$$\Longleftrightarrow$ $X_{n}^{\prime}$ $\geq 1-c_{n}$ (5.26)
@@
$P(F_{D_{n}}(D_{n},U_{n})\geq 1-x\mid A)$ $=\int_{0}^{1}P(F_{D_{n}}(D_{n},U_{n})\geq 1-x\mid A,U_{n}=u)\mathrm{\,d}u$ (5.27)
$=\int_{0}^{1}P(F_{D_{n}}(D_{n},u)\geq 1-x\mid A)\mathrm{\,d}u$
$=\frac{x}{c_{n}}P(D_{n}=1\mid A)\,,$
@@
$\int_{0}^{1}P(F_{D_{n}}(D_{n},u)\geq 1-x\mid A)\mathrm{\,d}u=\frac{x}{c_{n}}P(D_{n}=1\mid A)$
@@
$\int_{0}^{1}P(F_{D_{n}}(D_{n},u)\geq 1-x\mid A)\mathrm{\,d}u=\int_{0}^{1}P(D_{n}=1\mid A)\mathrm{\,d}u=\frac{x}{c_{n}}P(D_{n}=1\mid A).$
@@
$P(F_{D_{n}}(D_{n},U_{n})\geq 1-x\mid A)$ $=\int_{\frac{1-x}{1-\pi_{n}}}^{1}P(D_{n}=0\mid A)\mathrm{\,d}u+\int_{1-\frac{x}{\pi_{n}}}^{1}P(D_{n}=1\mid A)\mathrm{\,d}u$ (5.28)
$=\left(1-\frac{1-x}{1-c_{n}}\right)(1-P(D_{n}=1\mid A))+\frac{x}{c_{n}}P(D_{n}=1\mid A)$
$=\frac{x-c_{n}}{1-c_{n}}+\left(\frac{x}{c_{n}}-\frac{x-c_{n}}{1-c_{n}}\right)P(D_{n}=1\mid A)\,.$ (5.29)
@@
=$=\begin{cases}\frac{x}{c_{n}}P(D_{n}=1\mid Z=F_{Z}^{-1}(1-y),V=v_{y})\,,&\text{if }x\leq c_{n}\,,\\ \frac{x-c_{n}}{1-c_{n}}+\left(\frac{x}{c_{n}}-\frac{x-c_{n}}{1-c_{n}}\right)P(D_{n}=1\mid Z=F_{Z}^{-1}(1-y),V=v_{y})\,,&\text{if }x>c_{n}\,,\\ \end{cases}$
=$=\begin{cases}\frac{x}{c_{n}}P(D_{n}=1\mid Z=F_{Z}^{-1}(1-y))\,,&\text{if }x\leq c_{n}\,,\\ \frac{x-c_{n}}{1-c_{n}}+\left(\frac{x}{c_{n}}-\frac{x-c_{n}}{1-c_{n}}\right)P(D_{n}=1\mid Z=F_{Z}^{-1}(1-y))\,,&\text{if }x>c_{n}\,,\\ \end{cases}$
for some $v_{y}\in\mathbb{R}$ depending only on $F_{Z}$ and $y$ such that $F_{Z}(Z-)+V(F_{Z}(Z)-F_{Z}(Z-))=1-y$ for $Z=F_{Z}^{-1}(1-y)\,.$ For the third equality we also apply (5.27) and (5.28). The last equality follows from independence of $V$ from $D_{n}$ and $Z\,.$ Since by condition (III) the conditional default probability is increasing in $z\,,$ we obtain from (5.30) that $P(X_{n}\leq x\mid Y=y)$ is decreasing in $y\,.$ Hence, $C_{X_{n},Y}(u,v)=\int_{0}^{v}P(X_{n}\leq x\mid Y=v)\mathrm{\,d}v$ is concave in $v$ for all $u\,,$ which implies that $C_{X_{n},Y}$ is SI.
@@
$P\big{(}\bigcap\{X_{n}\leq x_{n}\}\mid Y=y\big{)}$ (5.31)
$=P\big{(}\bigcap\{X_{n}^{\prime}\geq 1-x_{n}\}\mid Y^{\prime}=1-y\big{)}$

### v2

@@ Acknowledgements
Appendix A Threshold Models
We represent siBBMs in terms of copula-based threshold models and provide a comparison result for threshold models that we use for the proof of Theorem 2.1.
A.1 Copula-Based Construction of siBMMs via Threshold Models
It is well known that every BMM can be represented as a threshold model, i.e., the default indicator variable $D_{n}$ is described by
$D_{n}=\mathds{1}_{\{X_{n}\leq c_{n}\}},\quad X_{n}=f_{n}(Y,\varepsilon_{n}),\quad n\in\{1,\ldots,N\},$ (A.1)
for some default threshold value $c_{n}$, some measurable function $f_{n}\colon\mathbb{R}^{2}\to\mathbb{R}$, and for independent idiosyncratic risk factors $\varepsilon_{1},\ldots,\varepsilon_{N}$ that are also independent from the random variable $Y$ which models the common risk factor; see, e.g., (McNeil et al., 2015, Section 8.4.4). If the distribution function of $X_{n}$ is continuous, then, under Assumption (I), it follows that
$c_{n}=F_{X_{n}}^{-1}(\pi_{n}).$ (A.2)
The dependence structure of $X_{n}$ and $Y$ is transferred to the dependence structure of $D_{n}$ and $Z,$ where we generally assume that $Z$ is a strictly decreasing transformation of $Y.$
$X_{n}=\sqrt{\rho_{n}}\,Y+\sqrt{1-\rho_{n}}\,\epsilon_{n}$ (A.3)
for i.i.d. standard normally distributed random variables $Y,\varepsilon_{1},\ldots,\varepsilon_{N}.$ In this case, $(X_{n},Y)$ follows a bivariate normal distribution with zero mean, variance $1,$ and factor loading $\sqrt{\rho_{n}},$ describing the sensitivity of borrower $n$ to the common factor $Y$. The asset return correlation between two distinct borrowers $n$ and $m$ is then given by $\sqrt{\rho_{n}\rho_{m}}$. The dependence structure between $X_{n}$ and $Y$ is described by a Gaussian copula with correlation $\sqrt{\rho_{n}}.$ It can easily be seen that $X_{n}$ in (A.3) is stochastically increasing in $Y$, and thus the implied BMM satisfies Assumption (III); see also Example A.5(a) below. However, in financial applications, the Gaussian copula is often not a good choice as it underestimates tail dependencies. This motivates to study robustness results for BMMs with general dependence structures.
In the sequel, we derive a copula-based characterization of siBMMs in terms of threshold models of the type (A.1). The following lemma shows that every threshold model $(D_{n})_{n}$ yields a siBMM if the function $f_{n}$ in (A.1) satisfies some monotonicity condition.
Lemma A.1 (Construction of siBMMs through threshold models).
Let $X_{n}=f_{n}(Y,\varepsilon_{n}),$ $n\in\{1,\ldots,N\},$ for some measurable function $f_{n}$ that is increasing in its first component and for $Y,\varepsilon_{1},\ldots,\varepsilon_{N}$ being independent. Consider the default indicator random variables $(D_{1},\ldots,D_{N})$ defined by $D_{n}:=\mathds{1}_{\{X_{n}\leq c_{n}\}}$ for some threshold values $c_{n}$ and define $Z:=g(Y)$ for some strictly decreasing function $g$. Then the following statements hold:
  1. (i)
  2. (ii)
    If additionally $Y,\varepsilon_{1},\ldots,\varepsilon_{N}$ have a continuous distribution function and if, for all $n\in\{1,\ldots,N\},$ $f_{n}$ is continuous, then also condition (I) is satisfied for $c_{n}=F_{X_{n}}^{-1}(\pi_{n}),$ i.e., $(D_{1},\ldots,D_{N},Z)$ is a siBMM with $\mathbb{P}(D_{n}=1)=\pi_{n}$ for all $n.$
We focus on the dependence structure of $(X_{n},Y),$ which we generally model with copulas. The aim will be to investigate the relation between the copula $C$ of $(X_{n},Y)$ and the function $f_{n}$ in Lemma A.1. To this end, recall that a bivariate copula is a distribution function on $0,1]^{2}$ with uniform univariate marginals. According to Sklar’s theorem (see e.g. Nelsen (2006)), every bivariate distribution function $F$ can be decomposed into its univariate marginal distribution functions $F_{1},F_{2}$ and a bivariate copula $C$ such that
$F(x_{1},x_{2})=C(F_{1}(x_{1}),F_{2}(x_{2}))\quad\text{for all }(x_{1},x_{2})\in[0,1]^{2}.$ (A.4)
The copula $C$ is uniquely determined on the Cartesian product $\mathsf{Ran}(F_{1})\times\mathsf{Ran}(F_{2}),$ where $\mathsf{Ran}(F_{i})$ denotes the range of $F_{i}.$ Further, for every bivariate copula $C$ and all univariate distribution functions $F_{1}$ and $F_{2},$ the right side of (A.4) defines a bivariate distribution function. Two important copulas are the product copula $\Pi(u,v):=uv,$ which models independence, and the upper Fréchet copula $M(u,v):=\min\{u,v\},$ which models comonotonicity. We refer to Nelsen (2006) and Durante and Sempi (2016) for an overview of copulas.
$C(u|v):=\partial_{2}C(u,v):=dC(u,v)/dv$
the partial derivative of $C$ with respect to its second component, which exists for Lebesgue-almost all $v;$ see e.g. (Durante and Sempi, 2016, Theorem 1.6.1). Then the conditional distribution function of $U|V=v$ can be represented, for fixed $u\in[0,1],$ by
$F_{U|V=v}(u)=\mathbb{P}(U\leq u\mid V=v)=C(u|v)$ (A.5)
$C^{-1}(t|v):=\inf\{u\in[0,1]\mid C(u|v)\geq t\}$ (A.6)
the generalized inverse of the function $u\mapsto C(u|v).$ For the construction of threshold models that imply BMMs satisfying monotonicity Assumption (III), we make use of the notion of stochastically increasing (SI) copulas, that are copulas of random vectors $(U,V)$ such that $U$ is SI in $V.$ As a consequence of (A.5), a bivariate copula $C$ is SI if and only if $C(u,v)$ is concave in $v$ for all $u.$ Many well known (sub-)families of copulas are SI such as the Gaussian copula with non-negative correlation parameter, the Clayton copula or the Gumbel-Hougaard copula; see Ansari and Rockel (2024) for an overview of various SI copula families from classes of Archimedean, extreme-value, and elliptical copulas.
Due to the following result, we can construct the function $f_{n}$ with the desired monotonicity condition in Lemma A.1 through a transformation of an SI copula that describes the dependence structure of $(X_{n},Y).$ Denote by $U(0,1)$ the uniform distribution on $(0,1)$.
Proposition A.2 (Copula-based construction of siBMMs through threshold models).
Let $C_{1},\ldots,C_{N}$ be SI copulas and let $Y,\varepsilon_{1},\ldots,\varepsilon_{n}\sim U(0,1)$ be i.i.d. random variables. For $f_{n}(v,t):=C_{n}^{-1}(t|v),$ with the notation defined in (A.6), consider $X_{n}:=f_{n}(Y,\varepsilon_{n}).$ Then the following statements hold true:
  1. (i)
    $X_{n}\sim U(0,1),$
  2. (ii)
    $f_{n}$ is componentwise increasing,
  3. (iii)
    $F_{X_{n},Y}=C_{X_{n},Y}=C_{n}.$
Remark A.3.
  1. (a)
    A general construction method for BMMs through threshold models is given in (McNeil et al., 2015, Lemma 11.10). Our Proposition A.2 is a copula-based version under the additional SI assumption.
  2. (b)
    A copula of the bivariate default vector $(D_{n},Z)$, constructed via the threshold model in Proposition A.2, is given by the survival copula of $C_{X_{n},Y}$, i.e.,
    $C_{D_{n},Z}(u,v)=\hat{C}_{X_{n},Y}(u,v):=u+v-1+C_{X_{n},Y}(1-u,1-v)$ (A.7)
    for $(u,v)\in[0,1]^{2}.$ This follows from the fact that both $D_{n}$ and $Z_{n}$ are decreasing transformations of $X_{n}$ and $Y,$ respectively; see e.g. (Nelsen, 2006, Theorem 2.4.4). [^5] The statement in this reference can be extended to random variables with discontinuous distribution function and to decreasing but not necessarily strictly decreasing transformations. Hence, for any copula of two random variables, its survival copula is a copula of the decreasing transformations. It can easily be verified that the survival copula $\hat{C}$ is SI if and only if $C$ is SI.
The following result is a reverse of Proposition A.2. It states that any siBMM can be written as a copula-based threshold model with SI copulas as in Proposition A.2. In other words, Proposition A.2 yields a general copula-based construction method of siBMMs.
Lemma A.4 (Representation of siBMMs through threshold models).
For any BMM $(D_{1},\ldots,D_{N},Z)$ satisfying conditions (I)–(III), there exist random variables $X_{1},\ldots,X_{N},Y\sim U(0,1)$ such that
  1. (i)
    $X_{1},\ldots,X_{N}$ are conditionally independent given $Y,$
  2. (ii)
    $C_{X_{n},Y}$ is SI, and
  3. (iii)
    $D_{n}=\mathds{1}_{\{X_{n}\leq\pi_{n}\}}$.
Further, a copula of $(D_{n},Z)$ is given by the survival copula $\hat{C}_{X_{n},Y}$ which is also SI.
In the following example, we discuss some well known SI copulas that can be used for the construction of siBMMs due to Proposition A.2.
Example A.5 (Copulas for siBMMs).
For modeling credit portfolio risks through threshold models of the form (A.1), lower tail dependencies between $X_{n}$ and the common risk factor $Y$ are crucial. Gaussian copulas, which are often used as a standard model, suffer from weak tail dependencies. Stronger lower tail dependencies between $X_{n}$ and $Y$ can be modeled, e.g., with Clayton copulas. The dependence structure of a Gaussian, a Clayton and a survival Clayton copula having Kendall’s tau value $\tau\approx 0.266$ is illustrated in Figure A.1.
  1. (a)
    The Gaussian copula $C_{\rho}^{Ga}(u,v)=\Phi_{\rho}(\Phi^{-1}(u),\Phi^{-1}(v))$ with correlation parameter $\rho\in[-1,1]$ is SI whenever $\rho\geq 0$ because $C_{\rho}^{Ga}(u|v)=\Phi((\Phi^{-1}(u)-\rho\,\Phi^{-1}(v))/\sqrt{1-\rho^{2}})$ is decreasing in $v$ for all $u.$ For a Gaussian copula model $F_{X_{n},Y}=C_{\rho}^{Ga},$ we have by Proposition A.2 the representation
    $X_{n}=f_{n}(Y,\varepsilon_{n})=(C_{\rho}^{Ga})^{-1}(\varepsilon_{n}|Y)=\Phi\left(\rho\,\Phi^{-1}(Y)+\sqrt{1-\rho^{2}}\,\Phi^{-1}(\varepsilon)\right),$ (A.8)
    which coincides with the single-factor Merton model in (A.3) for $\rho=\sqrt{\rho}_{n}$ up to a quantile transformation of $X_{n}$, $Y$, and $\varepsilon_{n}.$ Since the bivariate normal distribution is radially symmetric, see (Nelsen, 2006, Example 2.15), it follows from (A.7) and (Nelsen, 2006, Theorem 2.7.3) that $C_{D_{n},Z}=\hat{C}_{\rho}^{Ga}=C_{\rho}^{Ga}.$
    Figure A.1: Samples from a Gaussian copula (left), a Clayton copula (mid) and a survival Clayton copula (right), where the Gaussian copula parameter is chosen as the square root $\sqrt{\rho_{n}}\approx 0.405$ of the asset correlation obtained from the IRB formula (4.2) for $\pi_{n}=0.02.$ The parameter $\theta$ for the Clayton and survival Clayton copula is chosen as $\theta\approx 0.723$ so that their Kendall’s tau value $\tau\approx 0.266$ is the same as for the specified Gaussian copula. The copulas model the dependencies between the default indicator $X_{n}$ and the latent factor variable $Y$ in the respective threshold model.
  2. (b)
    The Clayton copula family $(C_{\theta}^{Cl})_{\theta\in[-1,\infty)}$ belongs to the class of Archimedean copulas and is given by $C_{\theta}^{Cl}(u,v)=\max\left\{0,\left(u^{-\theta}+v^{-\theta}-1\right)^{-1/\theta}\right\}$ for $\theta\neq 0$ and by $C_{\theta}^{Cl}(u,v)=uv$ for $\theta=0,$ see (Nelsen, 2006, Chapter 4) or (McNeil et al., 2015, Example 11.13). For $\theta>0,$ it holds that $C_{\theta}^{Cl}(u|v)=v^{-\theta-1}(u^{-\theta}+v^{-\theta}-1)^{-1/\theta-1},$ which decreases in $v,$ so that $C_{\theta}^{Cl}$ is SI. Taking the inverse of $C_{\theta}^{Cl}(u|v)$ with respect to $u,$ we obtain from Proposition A.2 for the Clayton copula model $F_{X_{n},Y}=C_{\theta}^{Cl}$ the closed-form expression
    $X_{n}=f_{n}(Y,\varepsilon_{n})=(C_{\theta}^{Cl})^{-1}(\varepsilon_{n}|Y)=\left[(\varepsilon_{n}Y^{\theta+1})^{-\theta/(\theta+1)}-Y^{-\theta}+1\right]^{-1/\theta}.$ (A.9)
    For $\theta>0$ the Clayton copula admits lower (but no upper) tail dependence; see e.g. (Ansari and Rockel, 2024, Table 3). Hence, the Clayton copula threshold model may be used for modeling default indicator variables where the underlying random variable $X_{n}$ depends for small values stronger on the risk factor $Y$ than in the Gaussian copula setting. From (A.7), we obtain that $C_{D_{n},Z}=\widehat{C_{\theta}^{Cl}}$.
  3. (c)
    $C_{\theta}^{sCl}(u,v):=\widehat{C_{\theta}^{Cl}}(u,v)=u+v-1+C_{\theta}^{Cl}(1-u,1-v)$ (A.10)
    for $(u,v)\in[0,1]^{2}.$ Due to straightforward symmetry arguments, we obtain for $\theta>0$ that $C_{\theta}^{sCl}$ is SI and exhibits upper but no lower tail dependence. The copulas of the siBMM, implied by the survival Clayton copula threshold model, are given by $C_{D_{n},Z}=\widehat{C_{\theta}^{sCl}}=C_{\theta}^{Cl}.$
A.2 Comparison of Threshold Models
For the proof of Theorem 2.1 and for the robustness results for siBMMs in terms of default integral functions, we use simple conditions on the bivariate SI copulas $C_{n}=C_{X_{n},Y},$ $n\in\{1\,\ldots,N\},$ that enable the comparison of losses in siBMMs relying on the copula-based threshold models in Proposition A.2.
The following proposition states that a pointwise comparison of SI copulas in threshold models yields a comparison of credit portfolio losses in convex order.
Proposition A.6 (Comparison of threshold models in convex order).
$\sum_{n=1}^{N}l_{n}D_{n}\leq_{cx}\sum_{n=1}^{N}l_{n}D_{n}^{\prime}.$ (A.11)
Example A.7.
Under the assumptions of Proposition A.6, assume that, for all $n,$ $(X_{n},Y)$ and $(X_{n}^{\prime},Y^{\prime})$ follow a Gaussian / Clayton / survival Clayton copula model, i.e., $C_{X_{n},Y}=C_{\rho_{n}}^{Ga}$ and $C_{X_{n}^{\prime},Y^{\prime}}=C_{\rho_{n}^{\prime}}^{Ga}$ for $\rho_{n},\rho_{n}^{\prime}\geq 0$ or $C_{X_{n},Y}=C_{\theta_{n}}^{Cl}$ and $C_{X_{n}^{\prime},Y^{\prime}}=C_{\theta_{n}^{\prime}}^{Cl}$ or $C_{X_{n},Y}=C_{\theta_{n}}^{sCl}$ and $C_{X_{n}^{\prime},Y^{\prime}}=C_{\theta_{n}^{\prime}}^{sCl}$ for $\theta_{n},\theta_{n}^{\prime}\geq 0.$ Since the family of Gaussian / Clayton / survival Clayton copulas is increasing in its parameter w.r.t. the pointwise order and the copulas are SI for non-negative parameters, we obtain that $\rho_{n}\leq\rho_{n}^{\prime}$, resp. $\theta_{n}\leq\theta_{n}^{\prime}$, for all $n$ implies (A.11). Further well known copula families that fulfill the assumptions of Proposition A.6 are given in Ansari and Rockel (2024).
Remark A.8.
If the distribution function of the common risk factor $Z$ is continuous [^6] If $F_{Z}$ is discontinuous, then the conditional default probability $\mathbb{P}(D_{n}=1\mid Z=F_{Z}^{-1}(s))$ can be similarly expressed as a generalized partial derivative of the underlying copula $C,$ see (Ansari and Rüschendorf, 2021, Theorem 2.2). Then the representation of $G_{D_{n},Z}$ in (A.12) has to be modified slightly., the default integral function $G_{D_{n},Z}$ has a simple representation through the copula of $(D_{n},Z)$ or $(X_{n},Y)$ as
$\begin{split}G_{D_{n},Z}(t)&=\int_{0}^{t}\mathbb{P}(D_{n}=1\mid Z=F_{Z}^{-1}(s))\mathrm{\,d}s=\int_{0}^{t}[1-\mathbb{P}(D_{n}=0\mid Z=F_{Z}^{-1}(s))]\mathrm{\,d}s\\ &=t-\int_{0}^{t}\partial_{2}C_{D_{n},Z}(1-\pi_{n},F_{Z}(F_{Z}^{-1}(s)))\mathrm{\,d}s=t-C_{D_{n},Z}(1-\pi_{n},t)\\ &=t-\hat{C}_{X_{n},Y}(1-\pi_{n},t)=\pi_{n}-C_{X_{n},Y}(\pi_{n},1-t),\end{split}$ (A.12)
where the first equality is due to the definition in (2.2). The third equality follows from the representation of the conditional distribution function in (A.5) and the transformation $V=F_{Z}(Z)$ using continuity of $F_{Z}.$ The fourth equality follows from $F_{Z}(F_{Z}^{-1}(s))=s$ for all $s$ using again continuity of $F_{Z}.$ The last two equalities are due to (A.7). Hence, by (A.12), a pointwise comparison of the default integral functions $G_{D_{n},Z}$ follows from a pointwise comparison of the underlying copulas in the siBMMs as in Proposition A.6.
Appendix B Proofs
@@
This appendix addresses the proofs of all results in this paper. We move the proof of Theorem 2.1 after the proof of Proposition A.6 because it is based on several results from Appendix A.1.
Proof of Lemma A.1:.
@@
$\mathbb{P}(D_{n}=1\mid Z=z)$ $=\mathbb{P}(X_{n}\leq c_{n}\mid Z=z)=\mathbb{P}(f_{n}(Y,\varepsilon_{n})\leq c_{n}\mid g(Y)=z)$
$=\mathbb{P}(f_{n}(g^{-1}(z),\varepsilon_{n})\leq c_{n}\mid Y=g^{-1}(z))=\mathbb{P}(f_{n}(g^{-1}(z),\varepsilon_{n})\leq c_{n}),$
@@
Under the continuity assumptions, it can easily be verified that $X_{n}$ has a continuous distribution function. Hence, $c_{n}=F_{X_{n}}^{-1}(\pi_{n})$ implies $\pi_{n}=\mathbb{P}(X_{n}\leq c_{n})=\mathbb{P}(D_{n}=1).$
Proof of Proposition A.2:.
@@
$F_{X_{n},Y}(x,y)$ $=\mathbb{P}(X_{n}\leq x,Y\leq y)=\int_{0}^{y}\mathbb{P}(f_{n}(Y,\varepsilon_{n})\leq x\mid Y=v)\mathrm{\,d}v$
$=\int_{0}^{y}\mathbb{P}(f_{n}(v,\varepsilon_{n})\leq x)\mathrm{\,d}v=\int_{0}^{y}\mathbb{P}(C_{n}^{-1}(\varepsilon_{n}|v)\leq x)\mathrm{\,d}v=\int_{0}^{y}\mathbb{P}(\varepsilon_{n}\leq C_{n}(x|v))\mathrm{\,d}v$
$=\int_{0}^{y}C_{n}(x|v)\mathrm{\,d}v=\int_{0}^{y}\partial C_{n}(x,v)/\partial v\mathrm{\,d}v=C_{n}(x,y)=C_{X_{n},Y}(x,y),$
where the third equality follows from independence of $Y$ and $\varepsilon_{n}.$ For the sixth equality, we use that $\varepsilon_{n}$ is uniformly distributed on $(0,1).$
@@
To show statement (ii), first observe that $f_{n}(v,t)$ is increasing in its second component $t$ because $C_{n}(t|v)$ and thus $C_{n}^{-1}(t|v)$ are increasing in $t.$ Further, $f_{n}(v,t)=C_{n}^{-1}(t|v)$ is increasing in $v$ for all tt\,. This follows from the fact that $C_{n}$ is SI, which in turn implies the concavity of $C_{n}(t,v)$ in $v$ for all $t$, and the fact that $C_{n}(t|v)=(\partial C_{n}(t,v)/\partial v)$ is decreasing in $v$ for all $t$ if and only if $C_{n}^{-1}(t|v)$ is increasing in $v$ for all $t$ (note that the inverse is w.r.t. the first component $t$ while the distribution functions are compared for different $v$).
Since $X_{n}\sim U(0,1),$ the default indicator $D_{n}$ satisfies condition (I). Properties (II) and (III) follow from Lemma A.1. ∎
Proof of Lemma A.4:.
We construct $X_{n}$ and $Y$ through the distributional transform of $D_{n}$ and $Z.$ To this end, define for a random variable $\xi$ its distributional transform by $\tau_{\xi}:=F_{\xi}(\xi,W):=F_{\xi}(\xi-)+W(F_{\xi}(\xi)-F_{\xi}(\xi-)),$ where $W\sim U(0,1)$ is independent from $\xi$ and where F(x)F(x-) denotes the left-hand limit of $F$ at $x.$ It is
$\tau_{\xi}$ $\sim U(0,1)\quad\text{and}\quad\xi=F_{\xi}^{-1}(\tau_{\xi})\quad P\text{-almost surely},$ (B.1)
@@
$X_{n}:=1-X_{n}^{\prime}\quad\text{and}\quad Y:=1-Y^{\prime}.$ (B.2)
Then (B.1) implies $X_{n},Y\sim U(0,1).$ Further, we have
$X_{n}$ $\leq c_{n}$ $\Longleftrightarrow$ $X_{n}^{\prime}$ $\geq 1-c_{n}$ (B.3)
@@
$\mathbb{P}(F_{D_{n}}(D_{n},U_{n})\geq 1-x\mid A)$ $=\int_{0}^{1}\mathbb{P}(F_{D_{n}}(D_{n},U_{n})\geq 1-x\mid A,U_{n}=u)\mathrm{\,d}u$ (B.4)
$=\int_{0}^{1}\mathbb{P}(F_{D_{n}}(D_{n},u)\geq 1-x\mid A)\mathrm{\,d}u=\frac{x}{c_{n}}\mathbb{P}(D_{n}=1\mid A),$
@@
$\int_{0}^{1}\mathbb{P}(F_{D_{n}}(D_{n},u)\geq 1-x\mid A)\mathrm{\,d}u=\frac{x}{c_{n}}\mathbb{P}(D_{n}=1\mid A)$
@@
$\int_{0}^{1}\mathbb{P}(F_{D_{n}}(D_{n},u)\geq 1-x\mid A)\mathrm{\,d}u=\int_{0}^{1}\mathbb{P}(D_{n}=1\mid A)\mathrm{\,d}u=\frac{x}{c_{n}}\mathbb{P}(D_{n}=1\mid A).$
@@
$\mathbb{P}(F_{D_{n}}(D_{n},U_{n})\geq 1-x\mid A)$ $=\int_{\frac{1-x}{1-\pi_{n}}}^{1}\mathbb{P}(D_{n}=0\mid A)\mathrm{\,d}u+\int_{1-\frac{x}{\pi_{n}}}^{1}\mathbb{P}(D_{n}=1\mid A)\mathrm{\,d}u$ (B.5)
$=\left(1-\frac{1-x}{1-c_{n}}\right)(1-\mathbb{P}(D_{n}=1\mid A))+\frac{x}{c_{n}}\mathbb{P}(D_{n}=1\mid A)$
$=\frac{x-c_{n}}{1-c_{n}}+\left(\frac{x}{c_{n}}-\frac{x-c_{n}}{1-c_{n}}\right)\mathbb{P}(D_{n}=1\mid A).$ (B.6)
@@
[…]
for some $v_{y}\in\mathbb{R}$ depending only on $F_{Z}$ and $y$ such that $F_{Z}(Z-)+V(F_{Z}(Z)-F_{Z}(Z-))=1-y$ for $Z=F_{Z}^{-1}(1-y).$ For the third equality we also apply (B.4) and (B.5). The last equality follows from independence of $V$ from $D_{n}$ and $Z.$ Since by condition (III) the conditional default probability is increasing in $z,$ we obtain from (B.7) that $\mathbb{P}(X_{n}\leq x\mid Y=y)$ is decreasing in $y.$ Hence, $C_{X_{n},Y}(u,v)=\int_{0}^{v}\mathbb{P}(X_{n}\leq u\mid Y=y)\mathrm{\,d}y$ is concave in $v$ for all $u,$ which implies that $C_{X_{n},Y}$ is SI.
@@
$P\big(\bigcap\{X_{n}\leq x_{n}\}\mid Y=y\big)=P\big(\bigcap\{X_{n}^{\prime}\geq 1-x_{n}\}\mid Y^{\prime}=1-y\big)$